### v2

# Poincaré inequality and Quantitative De Giorgi method for hypoelliptic operators

Abstract.
We propose a systematic elementary approach based on trajectories to prove Poincaré inequalities for hypoelliptic equations with an arbitrary number of Hörmander commutators, both in the local and in the non-local case. Our method generalises and simplifies the method introduced in [8] in the local case with one commutator, and extended in [13] to the non-local case with one commutator. It draws inspiration from the paper [17], although we use different trajectories. We deduce Harnack inequalities and Hölder regularity along the line of the De Giorgi method. Our results recover those in [2] in the local case, and are new in the non-local case.
AL is supported by the Cambridge Trust. CM is supported by ERC grant MAFRAN
1. Introduction
1.1. The question at hand
This paper is devoted to the regularity theory of hypoelliptic equations with rough coefficients. We consider equations with an arbitrary number of commutators to obtain regularity in all directions, in the sense of Hörmander [9], and whose diffusive part is either local (second-order) or non-local (fractional order). Beyond their structural intrinsic importance, such equations naturally appear in financial models (see Section 1.5).
We consider an unknown function $f=f(x,t)$ satisfying equations of the form
(1.1) $\partial_{t}f+{\mathcal{B}}f+{\mathcal{A}}^{*}{\mathcal{A}}f=0$
where ${\mathcal{B}}$ is a first-order differential operator and ${\mathcal{A}}$ is a differential, or fractional differential operator in some of the directions $x^{(0)}$ of $x$. We assume no regularity on the coefficients of the diffusive part ${\mathcal{A}}$. Then provided that ${\mathcal{A}}$ and ${\mathcal{B}}$ satisfy the Hörmander commutator condition [9], one expects the following Poincaré inequality to hold for weak non-negative sub-solutions:
$\int_{Q_{+}}\left(f-\frac{1}{|Q_{-}|}\int_{Q_{-}}f\right)_{+}\lesssim\int_{\Omega}|{\mathcal{A}}f|$
where $Q^{-},Q^{+}$ are two disjoint cylinders in the open domain $\Omega$ so that $Q^{+}$ is a time-translation of $Q^{-}$ into the future. In this paper we prove such an inequality with a quantitative control on the constant for a large class of equations of the type (1.1). The main novelty is the method itself, but the result is also new when $\mathcal{A}$ is non-local and more than one commutator is involved.
The proof is based on the combination of a method introduced in [8] and the explicit construction of appropriate new trajectories inspired partly from [17]. The Poincaré inequality is a powerful integral way to measure the oscillation of the sub-solution, and its use in this context goes back to [16, 15]. We deduce quantitatively the De Giorgi lemmas, the Hölder regularity for weak solutions, and the (weak) Harnack inequality for sub-solutions following the methodology recently developed in [8, 13].
1.2. Examples of equations
When $\kappa=1$, the prototypical local equation is the so-called Kolmogorov equation
(1.2) $\partial_{t}f+v\cdot\nabla_{x}f=\nabla_{v}\cdot\left[A\nabla_{v}f\right]$
on the unknown $f=f(x,v,t)$ depending on $x,v\in{\mathbb{R}}^{d}$ with $d\in{\mathbb{N}}$, $t\in{\mathbb{R}}$, where the matrix $A=A(x,v,t)$ is measurable symmetric so that, for some $\Lambda\geq 1$,
(1.3) $\forall\,(x,v,t)\in{\mathbb{R}}^{2d+1},\quad\Lambda^{-1}\leq A(x,v,t)\leq\Lambda.$
When the number of commutators $\kappa\geq 2$, the prototypical local equation is
(1.4) $\partial_{t}f+x^{(\kappa-1)}\cdot\nabla_{x^{(\kappa)}}f+x^{(\kappa-2)}\cdot\nabla_{x^{(\kappa-1)}}f+\dots+v\cdot\nabla_{x^{(1)}}f=\nabla_{v}\cdot\left[A\nabla_{v}f\right]$
on the unknown $f=f(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)$, depending on $x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v\in{\mathbb{R}}^{d}$, $t\in{\mathbb{R}}$, where the matrix $A$ is measurable, symmetric and so that, for some $\Lambda\geq 1$,
(1.5) $\forall\,(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)$ $\in{\mathbb{R}}^{(\kappa+1)d+1},$
$\Lambda^{-1}\leq A(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)\leq\Lambda.$
We denote $v=x^{(0)}$ for notational consistency.
In the fractional (non-local) case and with $\kappa=1$ commutators, the prototypical equation is the following fractional Kolmogorov equation
(1.6) $\partial_{t}f+v\cdot\nabla_{x}f=\left(-\Delta_{v}\right)^{\frac{\beta}{2}}\cdot\left[a\left(-\Delta_{v}\right)^{\frac{\beta}{2}}f\right]$
on the unknown $f=f(x,v,t)$ depending on $x,v\in{\mathbb{R}}^{d}$ and $t\in{\mathbb{R}}$, with $\beta\in(0,1)$ and a scalar function $a$ that is measurable and so that, for some $\Lambda\geq 1$,
(1.7) $a(x,v,t)\in[\Lambda^{-1},\Lambda].$
When the number of commutators $\kappa\geq 2$, the prototypical non-local equation is
(1.8) $\partial_{t}f+x^{(\kappa-1)}\cdot\nabla_{x^{(\kappa)}}f+x^{(\kappa-2)}\cdot\nabla_{x^{(\kappa-1)}}f+\dots+v\cdot\nabla_{x^{(1)}}f=\left(-\Delta_{v}\right)^{\frac{\beta}{2}}\cdot\left[a\left(-\Delta_{v}\right)^{\frac{\beta}{2}}f\right]$
on the unknown $f$ which depends on $x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},x^{(0)}\in{\mathbb{R}}^{d}$, $t\in{\mathbb{R}}$, and where the function $a$ is measurable and so that, for some $\Lambda\geq 1$,
(1.9) $\forall\,(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)\in{\mathbb{R}}^{(\kappa+1)d+1},\quad\Lambda^{-1}\leq a(x^{(\kappa)},x^{(\kappa-1)},\dots,x^{(1)},v,t)\leq\Lambda.$
In fact, our results still apply when the fractional diffusion operator on the right hand side of (1.6) and (1.8) is replaced by the more general nonlocal operator
$\mathcal{L}f\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t\right)$
$=\int_{{\mathbb{R}}^{d}}\left[f\left(x^{(\kappa)},\dots,x^{(1)},w,t\right)-f\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t\right)\right]K\left(x^{(\kappa)},\dots,x^{(1)},x^{(0)},w,t\right){\,\mathrm{d}}w$
where the kernel $K$ satisfies pointwise bounds
(1.10) $\forall\,\left(x^{(\kappa)},\dots,x^{(1)},v,w,t\right)\in{\mathbb{R}}^{(\kappa+2)d+1},\\ \Lambda^{-1}\big{|}v-w\big{|}^{-(d+2\beta)}\leq K\left(x^{(\kappa)},\dots,x^{(1)},v,w,t\right)\leq\Lambda\big{|}v-w\big{|}^{-(d+2\beta)},$
and the symmetry
(1.11) $\forall\,\left(x^{(\kappa)},\dots,x^{(1)},v,w,t\right)\in{\mathbb{R}}^{(\kappa+2)d+1},\\ K\left(x^{(\kappa)},\dots,x^{(1)},v,w,t\right)=K\left(x^{(\kappa)},\dots,x^{(1)},w,v,t\right).$
1.3. General assumptions (H)
We denote the variable $z:=(x,t)\in{\mathbb{R}}^{N+1}$ with $t\in{\mathbb{R}}$ and $x=(x^{(\kappa)},\dots,x^{(1)},x^{(0)})$ with each $x^{(i)}\in{\mathbb{R}}^{d_{i}}$, $d_{i}\in{\mathbb{N}}^{*}$ and $i\in\{0,\dots,\kappa\}$. We assume that $d_{0}\geq d_{1}\geq\dots\geq d_{\kappa}\geq 1$ and denote $N:=d_{0}+\cdots+d_{\kappa}$.
We consider (1.1) with operators ${\mathcal{A}}$ and ${\mathcal{B}}$ as follows. The operator ${\mathcal{B}}$ is given by
${\mathcal{B}}:=({\mathsf{B}}x)\cdot\nabla_{x},$
where the $N\times N$ matrix ${\mathsf{B}}$ has the form ($\mathbb{O}$ denotes a block matrix of zeros)
(1.12) ${\mathsf{B}}:=\begin{pmatrix}\mathbb{O}&{\mathsf{B}}_{\kappa}&\mathbb{O}&\cdots&\cdots&\mathbb{O}\\ \vdots&\mathbb{O}&{\mathsf{B}}_{\kappa-1}&\mathbb{O}&\cdots&\mathbb{O}\\ \vdots&\vdots&\mathbb{O}&\ddots&\mathbb{O}&\mathbb{O}\\ \vdots&\vdots&\vdots&\ddots&{\mathsf{B}}_{2}&\mathbb{O}\\ \vdots&\vdots&\vdots&\vdots&\mathbb{O}&{\mathsf{B}}_{1}\\ \mathbb{O}&\mathbb{O}&\cdots&\cdots&\cdots&\mathbb{O}\end{pmatrix},$
with $d_{i}\times d_{i-1}$ blocks ${\mathsf{B}}_{i}$ which are constant matrices of rank $d_{i}$.
The operator ${\mathcal{A}}$ is a linear operator acting only on the variable $x^{(0)}$ and $\beta$-homogeneous in this variable with $\beta\in(0,1]$ and it satisfies the following pointwise control from above and integral control from below, given $\varphi\in C^{\infty}_{c}({\mathbb{R}}^{N})$:
(1.13) $\|{\mathcal{A}}\varphi\|_{\infty}\lesssim\|\nabla_{(0)}\varphi\|_{L^{\infty}}^{\beta}\|\varphi\|_{L^{\infty}}^{1-\beta},\quad\int\left|{\mathcal{A}}\varphi\right|{\,\mathrm{d}}z\gtrsim\int\left|(-\Delta_{v})^{\beta/2}\varphi\right|{\,\mathrm{d}}z.$
The transport operator $\mathcal{T}:=\partial_{t}+{\mathcal{B}}$ naturally defines the non-commutative group of hyper-Galilean transformations, that generalises the Galilean transformations:
(1.14) $(\tilde{x},\tilde{t})\circ(x,t)=(x+\exp(t{\mathsf{B}})\tilde{x},t+\tilde{t}),\hskip 14.22636pt(x,t),\ (\tilde{x},\tilde{t})\in{\mathbb{R}}^{N+1}.$
This transformation leaves the operator $\mathcal{T}$ invariant: $(\mathcal{T}f)(\tilde{z}\circ z)=\mathcal{T}[f(\tilde{z}\circ\cdot)](z)$.
The scaling properties of ${\mathcal{A}}$ and ${\mathcal{B}}$ naturally define the hypoelliptic dilations for $r>0$:
(1.15) $\delta_{r}:=\textrm{diag}\left(r^{1+2\kappa\beta}\mathbb{I}_{d_{\kappa}},\dots,r^{1+2\beta}\mathbb{I}_{d_{1}},r\mathbb{I}_{d_{0}},r^{2\beta}\right)$
and we have the relation
$({\mathcal{T}}+{\mathcal{A}}^{*}{\mathcal{A}})[f\circ\delta_{r}]=r^{2\beta}\big{[}\left({\mathcal{T}}+{\mathcal{A}}^{*}{\mathcal{A}}\right)f\big{]}\circ\delta_{r}.$
These transformations allow to define the hypoelliptic cylinders
$Q_{r}(\tilde{z}):=\tilde{z}\circ\left(\delta_{r}\left(Q_{1}\right)\right)\quad\text{ with }\quad Q_{1}:=B_{1}\times\ldots\times B_{1}\times(-1,0]\subset{\mathbb{R}}^{N+1}.$
The (local) Kolmogorov equation (1.2) corresponds to $\kappa=1$, $d_{0}=d_{1}=d$, $(x^{(1)},x^{(0)})=(x,v)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$, ${\mathcal{A}}:=\sqrt{A}\nabla_{v}$ with a $d\times d$-matrix $A=A(x,v,t)$ measurable symmetric so that $A\in[\Lambda^{-1},\Lambda]$, and the matrix
(1.16) ${\mathsf{B}}:=\begin{pmatrix}\mathbb{O}&\mathbb{I}_{d}\\ \mathbb{O}&\mathbb{O}\end{pmatrix}.$
The (local) higher-order Kolmogorov equation (1.4) corresponds to $d_{i}=d$ for all $i$, $(x^{(\kappa)},\dots,x^{(1)},x^{(0)})\in{\mathbb{R}}^{(\kappa+1)d}$, ${\mathcal{A}}:=\sqrt{A}\nabla_{(0)}$ with a $d\times d$-matrix $A=A(x^{(\kappa)},\dots,x^{(1)},x^{(0)},t)$ measurable symmetric so that $A\in[\Lambda^{-1},\Lambda]$, and the matrix
${\mathsf{B}}:=\begin{pmatrix}\mathbb{O}&\mathbb{I}_{d}&\mathbb{O}&\ldots&\mathbb{O}\\ \mathbb{O}&\mathbb{O}&\mathbb{I}_{d}&\mathbb{O}&\vdots\\ \mathbb{O}&\vdots&\ddots&\ddots&\mathbb{O}\\ \vdots&\vdots&\vdots&\ddots&\mathbb{I}_{d}\\ \mathbb{O}&\mathbb{O}&\ldots&\ldots&\mathbb{O}\end{pmatrix}.$
The fractional Kolmogorov equations (1.6) and (1.8) are obtained along the same line with ${\mathcal{A}}:=\sqrt{a}(-\Delta_{v})^{\beta/2}$ and the same matrices ${\mathsf{B}}$.
1.4. Main result
Let us first give the definitions of weak sub-solutions and sub/super-solutions used in the results.
Definition 1.
Let $\Omega=\Omega_{x}\times\Omega_{t}=\Omega_{(N-1)}\times\cdots\times\Omega_{(1)}\times\Omega_{(0)}\times\Omega_{t}$ be an open subset of $\mathbb{R}^{N+1}$. A function $f$ is a locally integrable weak sub-solution of (1.1) if $f$ and ${\mathcal{A}}f$ are in $L^{1}_{loc}(\Omega)$ and satisfies
$\partial_{t}f+{\mathcal{B}}f+{\mathcal{A}}^{*}{\mathcal{A}}f\leq 0,$
in the distributional sense for non-negative test functions $\varphi\in C^{\infty}_{c}(\Omega)$. A function $f$ is a sub-solution of (1.1) if $f$ is a weak sub-solution in $L^{\infty}(\Omega_{t};L^{2}(\Omega_{x}))\cap L^{2}(\Omega_{(N-1)}\times\cdots\times\Omega_{(1)}\times\Omega_{t};H^{\beta}(\Omega_{(0)}))$ and satisfies
(1.17) $\partial_{t}f+{\mathcal{B}}f\in L^{2}(\Omega_{(N-1)}\times\cdots\times\Omega_{(1)}\times\Omega_{t};H^{-\beta}(\Omega_{(0)})).$
A function $f$ is a super-solution of (1.1) if $-f$ is a sub-solution.
Remark 2.
The condition (1.17) is used in [7] and it is weaker than the condition
$\partial_{t}f+{\mathcal{B}}f\in L^{2}(\Omega),$
used in [19]. In the definition of sub-solution, one could also consider a third condition, only valid for the local case, that reads as follows:
[…]
This condition was introduced in [8] and is weaker than the other two listed in the definition above. Indeed, it allows for instance to consider $f=f(t)=\mathbf{1}_{t\leq 0}$ as a sub-solution which explains the fact that the cylinders are disjoint in the following theorems.
Our first result is the Poincaré inequality:
Theorem 3.
Consider operators ${\mathcal{A}},{\mathcal{B}}$ satisfying the assumptions (H). Let $R>0$ be sufficiently large, depending on the number of commutators $\kappa$. Let $\Omega=B_{R}\times\dots\times B_{R}\times(-5,0)$ and suppose $f$ is a weak non-negative sub-solution to (1.1) in $\Omega$. Then
(1.18) $\int_{Q^{+}}\left(f-\frac{1}{|Q^{-}|}\int_{Q^{-}}f\right)_{+}{\,\mathrm{d}}z\lesssim\int_{\tilde{\Omega}}\big{|}{\mathcal{A}}f\big{|}{\,\mathrm{d}}z,$
where $Q^{+}:=Q_{1}\subset\Omega$ denotes the future cylinder, and $Q^{-}=B_{1}\times\ldots\times B_{1}\times(-5,-4]\subset\Omega$ denotes the past cylinder (see Figure 1). Moreover, if $\beta=1$, that is in the local case, $\tilde{\Omega}=\Omega$, whereas, if $\beta\in(0,1)$, that is in the non-local case, $\tilde{\Omega}={\mathbb{R}}^{d_{0}}\times B_{R}\times\dots\times B_{R}\times(-5,0)$. The constant is universal and depends on $\kappa$, $\beta$, $d_{0},\dots,d_{\kappa}$, ${\mathsf{B}}$, $R$ and $\Lambda$.
[…]
Figure 1. The different cylinders in the statement in Poincaré inequality (the time variable is represented vertically and upward).
Remarks 4.
  1. (1)
    The radius $R$ of the ambient space $Q_{R}$ is large enough such that the trajectories constructed in the proof (see (2.2)) do not exit $Q_{R}$. Since all maps are affine, we know that $R<+\infty$. One could try to optimise the size of the ambient space by optimising the control function used in the construction of the trajectories.
  2. (2)
    It is easy to include a source term $S$ in the right hand side of (1.1). Then the $L^{1}$ norm of $S$ on $\Omega$ should appear on the right hand side of (1.18).
  3. (3)
    It is likely that the approach can be extended to the case where the matrix ${\mathsf{B}}$ has non-zero terms above the sub-diagonal, since they do not break the commutator structure. This, however, would require additional techniques, which we leave for further work.
  4. (4)
    Under the same assumptions, it is straightforward to prove $L^{p}$ versions of the Poincaré inequality (1.18) by the same method, for $p\in(1,+\infty)$:
    $\int_{Q^{+}}\big{(}f-\langle f\rangle_{Q^{-}}\big{)}_{+}^{p}{\,\mathrm{d}}z\lesssim\int_{\Omega}\big{|}{\mathcal{A}}f\big{|}^{p}{\,\mathrm{d}}z.$
    In fact, it is even possible to obtain directly a gain of integrability for $f$ inside the Poincaré by our proof, see the related comments in Remark 8.
  5. (5)
    The left hand side of (1.18) could be mollified as follows
    (1.19) $\int_{\Omega}\left(f-\int_{\Omega}f\psi_{-}\right)\psi_{+}{\,\mathrm{d}}z\lesssim_{\psi_{\pm}}\int_{\Omega}\big{|}{\mathcal{A}}f\big{|}{\,\mathrm{d}}z$
    where $\psi_{\pm}\in C^{\infty}_{c}(Q^{\pm})$ are non-negative and integrate to one. This amounts to replacing respectively the indicator functions $\chi_{f(\cdot)\geq\langle f\rangle_{Q^{-}}}\chi_{Q^{+}}$ and $\chi_{Q^{-}}$ by some mollified versions $\psi_{+}$ and $\psi_{-}$ in the proof.
  6. (6)
    In such a mollified formulation, the Poincaré inequality in the local case would follow from the following “wholespace hypoelliptic Bogovoskiǐ result” (in the spirit of [4]): given $g=\psi_{+}-\psi_{-}$ with $\psi_{\pm}$ as above, construct a vector field $\mathbf{F}:=(F_{T},F_{(0)})$ with $F_{T}$ one-dimensional and $F_{(0)}$ $d_{0}$-dimensional, so that
    ${\mathcal{T}}^{*}F_{T}-\nabla_{(0)}\cdot F_{(0)}=g\quad\text{ and }\quad F_{T}\geq 0\quad\text{ and }\quad\|\nabla_{(0)}F_{T}\|_{\infty}+\|F_{(0)}\|_{\infty}\lesssim_{g}1.$
    This is an open question, but we believe this vector field to exist.
  7. (7)
    It is also possible to prove the local case of our Poincaré inequality through a “parabolic wholespace Bogovoskiǐ inequality” in $(x^{(\kappa)},\dots,x^{(1)},t)$ (excluding the $x^{(0)}$ variable, i.e. the velocity in case of a Kolmogorov equation), by following the arguments in [6]. This parabolic wholespace Bogovoskiǐ inequality is an easy generalisation of the Bogovoskiǐ inequality, where the explicit representation of Bogovoskiǐ is used to show that the first coordinate of the vector field solving the divergence problem is controlled below, due to the form of $g$ given in the previous point (6).
Our second main result is the weak Harnack inequality
Theorem 5 (Weak Harnack Inequality).
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
  • (i)
    ${\mathcal{A}}:=\sqrt{A}\nabla_{(0)}$ with $A\in[\Lambda^{-1},\Lambda]$ measurable symmetric matrix (local case),
  • (ii)
    ${\mathcal{A}}=\sqrt{a}(-\Delta_{v})^{\beta/2}$ with $a\in[\Lambda^{-1},\Lambda]$ scalar measurable (non-local case).
Let $f$ be a super-solution to (1.1) on $Q_{1}$ with $0\leq f\leq 1$ in $Q_{1}$ (local case) or on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-1,0]$ with $0\leq f\leq 1$ (non-local case). Then there is $C>0$ and $\zeta>0$ depending on $\beta,N,\kappa,\Lambda$ such that for $r_{0}>0$ sufficiently small the Weak Harnack inequality is satisfied:
(1.20) $\Bigg{(}\int_{\tilde{Q}^{-}_{r_{0}/2}}f^{\zeta}(z){\,\mathrm{d}}z\Bigg{)}^{\frac{1}{\zeta}}\leq C\inf_{Q_{r_{0}/2}}f,$
where $\tilde{Q}_{\frac{r_{0}}{2}}^{-}:=Q_{\frac{r_{0}}{2}}\big{(}(0,\dots,0,-\frac{5}{2}r_{0}^{2\beta}+\frac{1}{2}\big{(}\frac{r_{0}}{2}\big{)}^{2\beta})\big{)}\in{\mathbb{R}}^{N+1}$ (the local case is $\beta=1$).
Remark 6.
  1. (1)
    In the local case the strong Harnack follows immediately from the weak Harnack inequality above and the gain of integrability (first lemma of De Giorgi). By contrast, the strong Harnack inequality does not follow as immediately in the non-local case, since the function values outside the domain affect the solution inside the domain. It is however possible to adapt the recent work [12, 14], which proves the strong Harnack inequality in the non-local case with one commutator, to our setting.
  2. (2)
    It is standard to deduce the Hölder regularity from the weak Harnack inequality. We do not repeat the argument, but it is identical for instance to those in [8, 13].
1.5. Motivation
Kolmogorov equations appear in the theory of stochastic processes: equation (1.2) was studied by Kolmogorov in 1934 [11] when $A$ is the identity matrix to understand the “time-integrated Brownian motion”. The equation is then the Kolmogorov forward equation for the process
${\rm d}V_{t}={\rm d}W_{t},\quad{\rm d}X_{t}=V_{t}{\,\mathrm{d}}t.$
Kolmogorov obtained explicit formula for its fundamental solution in [11] which inspired the seminal work [9]. Integrating twice in time the Brownian motion would lead to equation (1.4) with $\kappa=2$ and $A$ the identity matrix. Due to this connection to stochastic processes, several mathematical models involving linear and nonlinear Kolmogorov type equations have also appeared in finance, and in particular equations of type (1.4) appear in various models for pricing of path-dependent financial instruments. For example the equation for $P=P(S,A,t)$
(1.21) $\partial_{t}P+\tfrac{1}{2}{\sigma}^{2}S^{2}\partial^{2}_{S}P+(\log S)\partial_{A}P+r(S\partial_{S}P-P)=0,\qquad S>0,\,A,t\in\mathbb{R},$
arises in the Black and Scholes option pricing problem
$\begin{cases}{\rm d}S_{t}=\mu S_{t}{\,\mathrm{d}}t+{\sigma}S_{t}{\,\mathrm{d}}W_{t},\\ {\rm d}A_{t}=S_{t}{\,\mathrm{d}}t,\end{cases}$
where ${\sigma}$ is the volatility of the stock price $S$, $\mu$ is the interest rate of a riskless bond and $P=P(S,A,t)$ is the price of the Asian option depending on the price of the stock $S$, the geometric average $A$ of the past price and the time to maturity $t$. For a more exhaustive treatment of the applications of Kolmogorov operators to finance and to stochastic theory, we refer to the monograph [18] by Pascucci.
Let us close this paragraph by mentioning that analogously, in the non-local case when $a=1$, or equivalently when $\mathcal{A}$ corresponds to the fractional Laplacian, the stochastic process underlying this fractional diffusion is a Lévy process.
1.6. Strategy of proof
The starting point is the method developed in [8] for proving the Poincaré inequality using trajectories and deducing the second Lemma of De Giorgi and the weak Harnack inequality.
The strategy to obtain a Poincaré inequality for the local case with one commutator in [8] can be summarised as follows. The first step is to mollify the characteristic function of $Q^{-}$, by adding an error term
$\int_{Q^{+}}\left(f-\langle f\rangle_{Q^{-}}\right)_{+}{\,\mathrm{d}}z\leq$ $\int_{z_{+}\in Q^{+}}\left(\int_{z_{-}\in Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(w,y){\,\mathrm{d}}z_{-}\right)_{+}{\,\mathrm{d}}z_{+}$
$+\varepsilon^{2d}\|f\|_{L^{2}(Q_{R})}$
with $0\leq\varphi_{\varepsilon}\leq 1$ smooth, $\varepsilon>0$, $z_{+}=(x,v,t)\in Q^{+}$ and $z_{-}=(y,w,s)\in Q^{-}$. The second step consists in constructing piecewise affine trajectories following the two vector fields $\mathcal{T}:=\partial_{t}+v\cdot\nabla_{x}$ and $\nabla_{v}$ to connect any $z_{+}\in Q^{+}$ with any $z_{-}\in Q^{-}$:
$(x,v,t)\underset{\nabla_{v}}{\longrightarrow}\left(x,\frac{x-y}{t-s},t\right)\underset{{\mathcal{T}}}{\longrightarrow}\left(y,\frac{x-y}{t-s},s\right)\underset{\nabla_{v}}{\longrightarrow}(y,w,s).$
Note here that a positive time gap between the two cylinders $Q^{-}$ and $Q^{+}$ is used to make sure the intermediate velocity $(x-y)/(t-s)$ remains bounded. We then write $f(z_{+})-f(z_{-})$ as an integral along the chosen trajectory and use ${\mathcal{T}}f+{\mathcal{A}}^{*}{\mathcal{A}}f\leq 0$:
$\int_{Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(z_{-}){\,\mathrm{d}}z_{-}$ $\lesssim\int_{Q^{-}}\int_{\text{trajectory}}{\mathcal{T}}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\lesssim\int_{Q^{-}}\int_{\text{trajectory}}\left[\nabla_{v}\cdot\left(A\nabla_{v}f\right)\right](\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}.$
We then want to integrate by parts the $v$-divergence in the past variable $z_{-}$. This integration by parts degenerates near the future point $z_{+}$, since $z_{+}$ does not depend on $z_{-}$, and it produces a non-integrable singularity. In order to overcome this singularity, the paper [8] introduced an additional small fourth sub-trajectory along $\nabla_{x}$:
$(v,x,t)$ $\underset{\nabla_{x}}{\longrightarrow}(v,x+\varepsilon w,t)\underset{\nabla_{v}}{\longrightarrow}\left(\frac{x+\varepsilon w-y}{t-s},x+\varepsilon w,t\right)$
$\underset{{\mathcal{T}}}{\longrightarrow}\left(\frac{x+\varepsilon w-y}{t-s},y,s\right)\underset{\nabla_{v}}{\longrightarrow}(w,y,s).$
This sub-trajectory “noises” the future position variable with the past velocity. This produces a second error term of the form $\varepsilon^{\sigma}\|f\|_{L^{1}_{t,v}W^{\sigma,1}_{x}}$, which can, however, be controlled for $\varepsilon$ small by the integral regularity $L^{1}_{t,v}W^{\sigma,1}_{x}$ (which can be established for sub-solutions to the Kolmogorov equation, for a small $\sigma>0$). Finally it yields
$\int_{Q^{+}}\big{(}f-\langle f\rangle_{Q^{-}}\big{)}_{+}{\,\mathrm{d}}z\lesssim C_{\varepsilon}\int_{Q_{R}}\big{|}S\big{|}{\,\mathrm{d}}z+\int_{Q_{R}}\big{|}\nabla_{v}f\big{|}{\,\mathrm{d}}z+\varepsilon^{\sigma}\|f\|_{L^{1}_{t,v}W^{\sigma,1}_{x}(Q_{R})}+\varepsilon^{2d}\|f\|_{L^{2}(Q_{R})}$
for a constant $C_{\varepsilon}>0$ depending on $\varepsilon>0$. The two additional errors weakens the Poincaré inequality, but the latter is shown to be sufficient for implementing the next steps of the De Giorgi theory in [8].
In the present article, we improve the method by simplifying the proof and removing the two error terms in the inequality. Moreover, the gain of integrability can directly be extracted from the proof of the Poincaré inequality. The first novel idea is to use three cylinders $Q^{-}$, $Q^{0}$ and $Q^{+}$, and mollify the characteristic function only in the intermediate cylinder $Q^{0}$, see Figure 2.
[…]
Figure 2. Construction of the trajectories. The curve $\Gamma_{+}$ connects any point $z_{+}\in Q^{+}$ to some intermediate point $z_{0}\in Q^{0}$, whereas the curve $\Gamma_{-}$ connects any point $z_{-}\in Q^{-}$ to some intermediate point $z_{0}\in Q^{0}$.
It is easy to deduce from the triangle inequality
(1.22) $\fint_{Q^{+}}\left[f(z_{+})-\langle f\rangle_{Q^{-}}\right]_{+}{\,\mathrm{d}}z_{+}$ $\leq\fint_{Q^{+}}\left(\fint_{Q^{0}}\left[f(z_{+})-f(z_{0})\right]\varphi_{\varepsilon}(v_{0},x_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{+}$
$\quad+\fint_{Q^{-}}\left(\fint_{Q^{0}}\left[f(z_{0})-f(z_{-})\right]\varphi_{\varepsilon}(v_{0},x_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{-}$
for $\varphi_{\varepsilon}\geq 0$ smooth, with mass $1$, where $z_{0}=(x_{0},v_{0},t_{0})\in Q^{0},z_{+}\in Q^{+},z_{-}\in Q^{-}$, and where we have denoted by $\fint_{Q}:=|Q|^{-1}\int_{Q}$ the normalized integral. We then construct trajectories to connect $z_{+}$ to $z_{0}$, and $z_{-}$ to $z_{0}$. This removes the first error term we had introduced in relation to the mollification $\varphi_{\varepsilon}$.
The second novel idea is to connect the points by a curved trajectory, rather than a piecewise affine one, by solving a control problem with a well-chosen forcing (i.e. control function), in order to reduce the order of the singularity of the integration by parts near $z_{\pm}$. More precisely, the first idea is to consider trajectories whose speed diverges at $z_{\pm}$ but with an integrable divergence so that the trajectories remain bounded. This idea is inspired from the reading of [17], but we propose simpler and more systematic trajectories. The second important idea is to use $\kappa+1$ linearly independent control power functions with such well-balanced diverging behaviour at $z_{\pm}$ in order to connect all points between the cylinders, thanks to the Hörmander commutator condition.
1.7. Outline
In Section 2 we provide the explicit construction of the trajectories connecting points in the future to points in the past. These trajectories are then used in Section 3 to prove the Poincaré inequality; they allow to estimate the $L^{1}$ norm in the future of the difference between $f$ and the past average of $f$, or in other words, the left hand side of (1.18). In Section 4, we prove the Harnack inequalities on the basis of this Poincaré inequality and deduce the Hölder regularity.
2. Construction of the trajectories
2.1. The base case
Let us first consider the case with one commutator ($\kappa=1$), $d_{1}=d_{0}=d$ and ${\mathsf{B}}_{1}=\text{Id}$. Given three points $z_{+}\in Q^{+},z_{0}\in Q^{0}$ and $z_{-}\in Q^{-}$, we want to construct two paths $s\rightarrow\Gamma_{+}(s)$ and $s\rightarrow\Gamma_{-}(s)$ for $s\in[0,1]$ such that (see Figures 2 and 3)
$\Gamma_{+}(s)=\left(X^{(1)}_{+}(s),X^{(0)}_{+}(s),T_{+}(s)\right)^{T},\qquad\Gamma_{+}(0)=z_{+},\qquad\Gamma_{+}(1)=z_{0},$
$\Gamma_{-}(s)=\left(X^{(1)}_{-}(s),X^{(0)}_{-}(s),T_{-}(s)\right)^{T},\qquad\Gamma_{-}(0)=z_{-},\qquad\Gamma_{-}(1)=z_{0},$
where
(2.1) $\begin{cases}{\frac{\mathrm{d}}{\mathrm{d}s}}X^{(0)}_{\pm}(s)=\mathsf{m}_{\pm}^{(0)}g_{0}^{\prime\prime}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}^{\prime\prime}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(1)}_{\pm}(s)=\delta_{\pm}X^{(0)}_{\pm}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}T_{\pm}(s)=\delta_{\pm},\end{cases}$
where $\delta_{\pm}:=t_{0}-t_{\pm}$ and the control functions $g_{i}\in C^{2}((0,1])$ with $g_{i}(0)=g_{i}^{\prime}(0)=0$ for $i=0,1$, and $\mathsf{m}^{(i)}_{\pm}\in{\mathbb{R}}^{d}$ for $i=0,1$.
-$\nabla_{(0)}$$\nabla_{(0)}$${\mathcal{T}}$$\dot{\Gamma}^{+}(s)$$x^{(0)}$$t$$x^{(1)}$$\Gamma_{+}(0)$$\Gamma_{+}(1)$
Figure 3. Construction of the trajectories. The curve $\Gamma_{+}$ connects any point $\Gamma_{+}(0)=z_{+}\in Q^{+}$ to some intermediate point $\Gamma_{+}(1)=z_{0}\in Q^{0}$ along the vector fields ${\mathcal{T}}$p and $\nabla_{(0)}$.
Solving the differential equations yields
(2.2) $\begin{cases}X^{(0)}_{\pm}(s)=x^{(0)}_{\pm}+\mathsf{m}_{\pm}^{(0)}g_{0}^{\prime}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}^{\prime}(s),\\[5.69054pt] X^{(1)}_{\pm}(s)=x^{(1)}_{\pm}+s\delta_{\pm}x^{(0)}_{\pm}+\delta_{\pm}\left[\mathsf{m}_{\pm}^{(0)}g_{0}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}(s)\right],\\[5.69054pt] T_{\pm}(s)=st_{0}+(1-s)t_{\pm}.\end{cases}$
Let us denote $\text{{\bf 1}}_{d}$ the $d$-vector of $1$’s, $\text{{\bf 0}}_{d}$ the $d$-vector of 00’s, $\text{Id}_{d}$ the $d\times d$ identity matrix, $\mathbb{O}$ a zero block of arbitrary size, and
$\mathsf{M}_{\pm}:=\begin{pmatrix}\mathsf{m}^{(0)}_{\pm}\\ \mathsf{m}^{(1)}_{\pm}\end{pmatrix},\quad{\mathsf{W}}(s):=\begin{pmatrix}g_{0}^{\prime}(s)\text{Id}_{d}&g_{1}^{\prime}(s)\text{Id}_{d}\\ g_{0}(s)\text{Id}_{d}&g_{1}(s)\text{Id}_{d}\end{pmatrix},$
${\mathsf{W}}_{\pm}^{\delta}(s):=\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \mathbb{O}&\delta_{\pm}\text{Id}_{d}\end{pmatrix}{\mathsf{W}}(s),\quad\mathsf{Y}_{\pm}=\begin{pmatrix}\mathsf{y}^{(0)}_{\pm}\\ \mathsf{y}^{(1)}_{\pm}\end{pmatrix}:=\begin{pmatrix}x^{(0)}_{0}-x^{(0)}_{\pm}\\ x^{(1)}_{0}-x^{(1)}_{\pm}-\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}.$
The boundary conditions $\Gamma_{\pm}(1)=z_{0}$ impose
(2.3) ${\mathsf{W}}_{\pm}^{\delta}(1)\mathsf{M}_{\pm}=\mathsf{Y}_{\pm}\quad\Longrightarrow\quad\mathsf{M}_{\pm}={\mathsf{W}}_{\pm}^{\delta}(1)^{-1}\mathsf{Y}_{\pm},$
provided that the Wronskian matrix ${\mathsf{W}}(1)$ is invertible at $s=1$. We deduce
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}={\mathsf{W}}_{\pm}^{\delta}(s)\mathsf{M}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}.$
Using (2.3) we obtain
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}$ $={\mathsf{W}}_{\pm}^{\delta}(s){\mathsf{W}}_{\pm}^{\delta}(1)^{-1}\mathsf{Y}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
$=W^{\delta}_{\pm}(s)\big{[}W^{\delta}_{\pm}(1)\big{]}^{-1}\begin{pmatrix}x_{0}^{(0)}\\ x_{0}^{(1)}\end{pmatrix}$
$\quad-W^{\delta}_{\pm}(s)\big{[}W^{\delta}_{\pm}(1)\big{]}^{-1}\begin{pmatrix}x^{(0)}_{\pm}\\ x^{(1)}_{\pm}+\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
$=\mathfrak{A}^{s}\begin{pmatrix}x_{0}^{(0)}\\ x_{0}^{(1)}\end{pmatrix}+\mathfrak{B}^{s}=\mathfrak{A}^{s}x_{0}+\mathfrak{B}^{s}=:\Phi^{s}_{\pm}(x_{0})$
is an affine function with matrix $\mathfrak{A}^{s}:={\mathsf{W}}_{\pm}^{\delta}(s)[{\mathsf{W}}_{\pm}^{\delta}(1)]^{-1}$ and a vector $\mathfrak{B}^{s}$ that depend only on $s$ and $x_{\pm}$. We still have to prove that the matrix $\mathfrak{A}^{s}$ is invertible for $s\in(0,1]$. If so, given $s\in(0,1]$, the derivative along the first variable $X^{(0)}_{\pm}(s)$ of the inverse is
$\nabla_{(0)}\left(\Phi_{\pm}^{s}\right)^{-1}={\mathsf{W}}^{\delta}_{\pm}(1)\big{[}{\mathsf{W}}^{\delta}_{\pm}(s)\big{]}^{-1}\left(\begin{matrix}\text{{\bf 1}}_{d}\\ \text{{\bf 0}}_{d}\end{matrix}\right).$
We choose $g_{i}(s):=(1+\alpha_{i})^{-1}(2+\alpha_{i})^{-1}s^{2+\alpha_{i}}$, $i=0,1$, with $\alpha_{0},\alpha_{1}\in(-1,0)$ and $\alpha_{0}\not=\alpha_{1}$. Then the Wronskian matrix is invertible for all $s\not=0$:
(2.4) $\det{\mathsf{W}}(s)$ $=\left(g_{0}^{\prime}(s)g_{1}(s)-g_{0}(s)g_{1}^{\prime}(s)\right)^{d}$
=$\frac{(\alpha_{0}-\alpha_{1})^{d}}{(1+\alpha_{0})^{d}(1+\alpha_{1})^{d}(2+\alpha_{0})^{d}(2+\alpha_{1})^{d}}s^{(3+\alpha_{0}+\alpha_{1})d}\not=0.$
Our choice of two linearly independent control functions $g_{0}$ and $g_{1}$ is guided by ensuring the Wronskian matrix to be invertible. Moreover,
${\mathsf{W}}_{\pm}^{\delta}(1):=\begin{pmatrix}\frac{\text{Id}_{d}}{(1+\alpha_{0})}&\frac{\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}\end{pmatrix},$
${\mathsf{W}}_{\pm}^{\delta}(s)^{-1}:=\frac{1}{\delta_{\pm}\det{\mathsf{W}}(s)}\begin{pmatrix}\frac{\delta_{\pm}s^{2+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}&\frac{-s^{1+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{-\delta_{\pm}s^{(2+\alpha_{0})}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{s^{1+\alpha_{0}}\text{Id}_{d}}{(1+\alpha_{0})}\end{pmatrix},$
which finally yields
$\nabla_{(0)}\left(\Phi_{\pm}^{s}\right)^{-1}$ $=W^{\delta}_{\pm}(1)\big{[}W^{\delta}_{\pm}(s)\big{]}^{-1}\begin{pmatrix}\text{{\bf 1}}_{d}\\ \text{{\bf 0}}_{d}\end{pmatrix}$
$=\frac{1}{(\alpha_{0}-\alpha_{1})}\begin{pmatrix}(2+\alpha_{0})s^{-1-\alpha_{0}}-(2+\alpha_{1})s^{-1-\alpha_{1}}\text{{\bf 1}}_{d}\\ \delta_{\pm}\left(s^{-1-\alpha_{0}}-s^{-1-\alpha_{1}}\right)\text{{\bf 1}}_{d}\end{pmatrix}$
$=O(s^{-1-\alpha_{0}})+O(s^{-1-\alpha_{1}}),$
which remains integrable for $\alpha_{0},\alpha_{1}\in(-1,0)$ with $\alpha_{0}\neq\alpha_{1}$. Observe also that our choice of control functions $g_{0}$ and $g_{1}$ implies that all their derivatives up to order two are integrable on $s\in[0,1]$, which implies that the trajectories are bounded (with integrable tangent vector field).
2.2. The general case
Let us now consider the general case with $\kappa\geq 2$ commutators, general dimensions $d_{0}\geq d_{1}\geq\dots\geq d_{\kappa}\geq 1$ and surjective $d_{i}\times d_{i-1}$ matrices ${\mathsf{B}}_{i}$. Consider for $s\in[0,1]$ the paths
$\Gamma_{+}(s)=\left(X^{(\kappa)}_{+}(s),\dots,X^{(0)}_{+}(s),T_{+}(s)\right)^{T},\qquad\Gamma_{+}(0)=z_{+},\qquad\Gamma_{+}(1)=z_{0},$
$\Gamma_{-}(s)=\left(X^{(\kappa)}_{-}(s),\dots,X^{(0)}_{-}(s),T_{-}(s)\right)^{T},\qquad\Gamma_{-}(0)=z_{-},\qquad\Gamma_{-}(1)=z_{0},$
and the following control problem
(2.5) $\begin{cases}{\frac{\mathrm{d}}{\mathrm{d}s}}X^{(0)}_{\pm}(s)=\sum_{i=0}^{\kappa}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa+1)}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(1)}_{\pm}(s)=\delta_{\pm}{\mathsf{B}}_{1}X^{(0)}_{\pm}(s),\\[5.69054pt] \quad\vdots\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(\kappa)}_{\pm}(s)=\delta_{\pm}{\mathsf{B}}_{\kappa}X^{(\kappa-1)}_{\pm}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}T_{\pm}(s)=\delta_{\pm}\end{cases}$
with $\delta_{\pm}:=t_{0}-t_{\pm}$ and the control functions $g_{i}\in C^{\kappa+1}((0,1])$ so that $g_{i}(0)=g_{i}^{\prime}(0)=\dots=g_{i}^{(\kappa)}(0)=0$ for $i=0,1,\dots,\kappa$, and $\mathsf{m}^{(i)}_{\pm}\in{\mathbb{R}}^{d_{0}}$ for $i=0,\dots,\kappa$. Then
(2.6) $\begin{cases}X^{(0)}_{\pm}(s)=x^{(0)}_{\pm}+\sum_{i=0}^{\kappa}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa)}(s),\\[5.69054pt] X^{(1)}_{\pm}(s)=x^{(1)}_{\pm}+s\delta_{\pm}\tilde{\mathsf{B}}_{1,1}x^{(0)}_{\pm}+\delta_{\pm}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{1,1}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa-1)}(s)\right],\\[5.69054pt] X^{(2)}_{\pm}(s)=x^{(2)}_{\pm}+s\delta_{\pm}\tilde{\mathsf{B}}_{2,2}x^{(1)}_{\pm}+\frac{(s\delta_{\pm})^{2}}{2}\tilde{\mathsf{B}}_{2,1}x^{(0)}_{\pm}+\delta_{\pm}^{2}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{2,1}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa-2)}(s)\right],\\[5.69054pt] \quad\vdots\\[5.69054pt] X^{(\kappa)}_{\pm}(s)=x_{\pm}^{(\kappa)}+\sum_{i=1}^{\kappa}\frac{(s\delta_{\pm})^{i}}{i!}\tilde{\mathsf{B}}_{\kappa,\kappa-i+1}x^{(\kappa-i)}_{\pm}+\delta_{\pm}^{\kappa}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{\kappa,1}\mathsf{m}_{\pm}^{(i)}g_{i}(s)\right],\\[5.69054pt] T_{\pm}(s)=st_{0}+(1-s)t_{\pm}\end{cases}$
where $\tilde{\mathsf{B}}_{i,j}:={\mathsf{B}}_{i}{\mathsf{B}}_{i-1}\cdots{\mathsf{B}}_{j}$ for $1\leq j\leq i$, which is a $d_{i}\times d_{j-1}$ block matrix. We denote
(2.7) $\mathsf{M}_{\pm}:=\begin{pmatrix}\mathsf{m}^{(0)}_{\pm}\\ \vdots\\ \mathsf{m}^{(\kappa)}_{\pm}\end{pmatrix},\quad{\mathsf{W}}(s):=\begin{pmatrix}g_{0}^{(\kappa)}(s)\text{Id}_{d_{0}}&\dots&g_{\kappa}^{(\kappa)}(s)\text{Id}_{d_{0}}\\ \vdots&\vdots&\vdots\\ g_{0}(s)\text{Id}_{d_{0}}&\dots&g_{\kappa}(s)\text{Id}_{d_{0}}\end{pmatrix},$
(2.8) ${\mathsf{W}}_{\pm}^{\delta}(s):=\begin{pmatrix}\text{Id}_{d_{0}}&\mathbb{O}&\cdots&\cdots&\mathbb{O}\\ \mathbb{O}&\tilde{\mathsf{B}}_{1,1}\delta_{\pm}&\mathbb{O}&\cdots&\mathbb{O}\\ \vdots&\mathbb{O}&\tilde{\mathsf{B}}_{2,1}\delta_{\pm}^{2}&\ddots&\mathbb{O}\\ \vdots&\vdots&\ddots&\ddots&\mathbb{O}\\ \mathbb{O}&\cdots&\cdots&\mathbb{O}&\tilde{\mathsf{B}}_{\kappa,1}\delta_{\pm}^{\kappa}\end{pmatrix}W(s)=:\mathsf{R}\,\mathsf{W}(s),$
(2.9) $\mathsf{Y}_{\pm}=\begin{pmatrix}\mathsf{y}^{(0)}_{\pm}\\ \vdots\\ \vdots\\ \mathsf{y}^{(\kappa)}_{\pm}\end{pmatrix}:=\begin{pmatrix}\mathsf{x}_{0}^{(0)}\\ \vdots\\ \vdots\\ \mathsf{x}_{0}^{(\kappa)}\end{pmatrix}-\mathsf{T}_{\pm}(1)\begin{pmatrix}\mathsf{x}^{(0)}_{\pm}\\ \vdots\\ \vdots\\ \mathsf{x}^{(\kappa)}_{\pm}\end{pmatrix},$
(2.10) ${\mathsf{T}}_{\pm}(s):=\begin{pmatrix}\text{Id}_{d_{0}}&\mathbb{O}&\cdots&\cdots&\mathbb{O}\\ (s\delta_{\pm})\tilde{\mathsf{B}}_{1,1}&\hbox{Id}_{d_{1}}&\mathbb{O}&\cdots&\mathbb{O}\\ \vdots&(s\delta_{\pm})\tilde{\mathsf{B}}_{2,2}&\hbox{Id}_{d_{2}}&\ddots&\mathbb{O}\\ \vdots&\vdots&\ddots&\ddots&\mathbb{O}\\ \frac{(s\delta_{\pm})^{\kappa}}{\kappa!}\tilde{\mathsf{B}}_{\kappa,1}&\cdots&\cdots&(s\delta_{\pm})\tilde{\mathsf{B}}_{\kappa,\kappa}&\hbox{Id}_{d_{\kappa}}\end{pmatrix}.$
Note that the matrix ${\mathsf{W}}(s)$ is the Wronskian of the family of functions $(g_{i})_{i=0}^{\kappa}$ and is invertible when they are the linearly independent solutions to a $(\kappa+1)$-order linear ODE. The boundary conditions $\Gamma_{\pm}(1)=z_{0}$ impose
${\mathsf{W}}_{\pm}^{\delta}(1)\mathsf{M}_{\pm}=\mathsf{Y}_{\pm}\quad\Longrightarrow\quad\mathsf{M}_{\pm}={\mathsf{W}}_{\pm}^{\delta}(1)^{-1}\mathsf{Y}_{\pm}.$
Note the pseudo-inverse of the matrix ${\mathsf{W}}_{\pm}^{\delta}(1)$ always exists, and is given by the product of the inverse of ${\mathsf{W}}(1)$ times the left inverse of $\mathsf{R}$. Thus we deduce that
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ \ \vdots\\ X^{(\kappa)}_{\pm}(s)\end{pmatrix}$ $={\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}x_{0}+\left({\mathsf{T}}_{\pm}(s)-{\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}{\mathsf{T}}_{\pm}(1)\right)x_{\pm}$
(2.11) $=\mathfrak{A}^{s}x_{0}+\mathfrak{B}^{s}=:\Phi^{s}_{\pm}(x_{0})$
is an affine function with matrix $\mathfrak{A}^{s}:={\mathsf{W}}_{\pm}^{\delta}(s)[{\mathsf{W}}_{\pm}^{\delta}(1)]^{-1}=\mathsf{R}\mathsf{W}(s)\mathsf{W}(1)^{-1}\mathsf{R}^{-1}$ and vector $\mathfrak{B}^{s}$ that depend only on $s$ and $x_{\pm}$. We still have to prove that the matrix $\mathfrak{A}^{s}$ is invertible for $s\in(0,1]$. If so, given $s\in(0,1]$, the derivative along the first variable $X^{(0)}_{\pm}(s)$ of the inverse is
(2.12) $\nabla_{(0)}\left(\Phi_{\pm}^{s}\right)^{-1}={\mathsf{W}}_{\pm}^{\delta}(1)\big{[}{\mathsf{W}}_{\pm}^{\delta}(s)\big{]}^{-1}\begin{pmatrix}\hbox{Id}_{d_{0}}\\ \mathbb{O}_{d_{1}}\\ \vdots\\ \mathbb{O}_{d_{\kappa}}\end{pmatrix}.$
We finally choose the control functions
(2.13) $g_{i}(s):=\frac{s^{1+\kappa+\alpha_{i}}}{(1+\alpha_{i})(2+\alpha_{i})\cdots(1+\kappa+\alpha_{i})}$
with $\alpha_{i}\in(-1,0)$ pairwise distinct for $i=0,\dots,\kappa$. Thus, we are left to prove that $\mathsf{W}(1)$ and $\mathfrak{A}_{s}$ are invertible for $s\in(0,1]$, but this boils down to prove the matrix $\mathsf{W}(s)$ is invertible for $s\in(0,1]$. Indeed, when $s\neq 0$ the Wronskian matrix is invertible with the precise $s$-behaviour given by the following lemma:
Lemma 7.
Consider the $(\kappa+1)\times(\kappa+1)$ matrix $P(s):=(s^{1+\alpha_{j}+i}/p_{i,j})_{i,j=0}^{\kappa}$ with $p_{i,j}:=(1+\alpha_{j})\cdots(1+i+\alpha_{j})$ the matrix [^1] The notation $P\otimes\text{Id}_{d_{0}}$ is the classical tensor product between the matrices $P$ and $\text{Id}_{d_{0}}$. $W:=P\otimes\text{Id}_{d_{0}}$. Then their determinants are given by
(2.14) $\det P(s)=\left(\frac{\prod_{i,j=0|i<j}^{\kappa}(\alpha_{i}-\alpha_{j})}{\prod_{i,j=0}^{\kappa}(1+i+\alpha_{j})}\right)s^{\frac{(\kappa+1)(\kappa+2)}{2}+\sum_{i=0}^{\kappa}\alpha_{i}}$
$\det{\mathsf{W}}(s)=\left(\frac{\prod_{i,j=0|i<j}^{\kappa}(\alpha_{i}-\alpha_{j})^{d_{0}}}{\prod_{i,j=0}^{\kappa}(1+i+\alpha_{j})^{d_{0}}}\right)s^{d_{0}\frac{(\kappa+1)(\kappa+2)}{2}+d_{0}\sum_{i=0}^{\kappa}\alpha_{i}}.$
Proof of Lemma 7.
The factorisation in $s$ gives the exponents. It remains to compute the determinant of matrix ${P}(1):=(1/p_{i,j})_{i,j=0}^{\kappa}$. From now on, for every $n\geq 1$, we define $Q_{n}(X)=\prod_{j=0}^{n-1}(X+\kappa+1-j)$ a polynomial of degree $n$. Then, we observe:
$\prod_{i,j=0}^{\kappa}(1+i+\alpha_{j})\det{P}(1)=\det\begin{pmatrix}Q_{\kappa}(\alpha_{0})&Q_{\kappa}(\alpha_{1})&\cdots&Q_{\kappa}(\alpha_{\kappa})\\ Q_{\kappa-1}(\alpha_{0})&Q_{\kappa-1}(\alpha_{1})&\cdots&Q_{\kappa-1}(\alpha_{\kappa})\\ \vdots&\ddots&\ddots&\vdots\\ Q_{1}(\alpha_{0})&Q_{1}(\alpha_{1})&\cdots&Q_{1}(\alpha_{\kappa})\\ 1&1&\cdots&1\end{pmatrix},$
where the equivalence follows by multiplying the $j^{th}$-column by $\prod\limits_{i=0}^{\kappa}(1+i+\alpha_{j})$ according to classical laws of multiplication by scalars for determinants of matrices. Then the determinant we are interested in is equal to the determinant of the equivalent matrix that comes from performing linear combinations of rows to get $\alpha_{j-1}^{\kappa+1-i}$ on the $i^{th}$ row and $j^{th}$ column which gives a Vandermonde determinant of value $\prod_{i,j=0|i<j}^{\kappa}(\alpha_{i}-\alpha_{j})$. The determinant of $W$ directly follows by tensor calculus. ∎
Then, by Lemma 7, we thus find
${\mathsf{W}}(s)^{-1}:=\frac{1}{\det{\mathsf{W}}(s)}\operatorname{Comatrix}{\mathsf{W}}(s).$
Going back to the calculation of (2.12), and combining (2.14) with the $s$-scaling of the entries of the cofactor matrix and our choice (2.13), we get
(2.15) $\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}={\mathsf{W}}_{\pm}^{\delta}(1)\big{[}{\mathsf{W}}_{\pm}^{\delta}(s)\big{]}^{-1}\begin{pmatrix}\hbox{Id}_{d_{0}}\\ \mathbb{O}_{d_{1}}\\ \vdots\\ \mathbb{O}_{d_{\kappa}}\end{pmatrix}=\sum_{i=0}^{\kappa}O(s^{-1-\alpha_{i}}),$
which is integrable for $\alpha_{i}\in(-1,0)$. Observe also that our choice of control functions $g_{i}$, $i=1,\dots,\kappa$, implies that all their derivatives up to order $\kappa+1$ are integrable on $s\in[0,1]$, which implies that the trajectories are bounded (with integrable tangent vector field).
3. Proof of the Poincaré inequality
We consider the three cylinders $Q^{+},Q^{-},Q^{0}\subset\Omega$ as in Figure 2, operators $\mathcal{A}$ and $\mathcal{B}$ that satisfy (H), and a subsolution $f$ of (1.1). Let $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{N})$ be a non-negative function in the first $N$ variables $(x^{(\kappa)},\dots,x^{(0)})$ (excluding the time component) with compact support in any time-slice of $Q^{0}$ and such that $\fint_{Q^{0}}\varphi{\,\mathrm{d}}x=1$. Then
(3.1) $\fint_{Q^{+}}\Big{(}f(z_{+})-\langle f\rangle_{Q^{-}}\Big{)}_{+}{\,\mathrm{d}}z_{+}$ $\leq\fint_{Q^{+}}\fint_{Q^{-}}\Big{(}f(z_{+})-f(z_{-})\Big{)}_{+}{\,\mathrm{d}}z_{-}{\,\mathrm{d}}z_{+}$
$\leq\fint_{Q^{+}}\fint_{Q^{-}}\Big{(}f(z_{+})-\langle f\varphi\rangle_{Q^{0}}\Big{)}_{+}{\,\mathrm{d}}z_{-}{\,\mathrm{d}}z_{+}$
$\quad+\fint_{Q^{+}}\fint_{Q^{-}}\Big{(}\langle f\varphi\rangle_{Q^{0}}-f(z_{-})\Big{)}_{+}{\,\mathrm{d}}z_{-}{\,\mathrm{d}}z_{+}$
$\leq\fint_{Q^{+}}\Bigg{\{}\underbrace{\fint_{Q^{0}}\big{(}f(z_{+})-f(z_{0})\big{)}\varphi(x_{0}){\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{+}}\Bigg{\}}_{+}{\,\mathrm{d}}z_{+}$
$\quad+\fint_{Q^{-}}\Bigg{\{}\underbrace{\fint_{Q^{0}}\big{(}f(z_{0})-f(z_{-})\big{)}\varphi(x_{0}){\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{-}}\Bigg{\}}_{+}{\,\mathrm{d}}z_{-}.$
where $z_{+}=(x_{+},t_{+}),z_{-}=(x_{-},t_{-})$ and $z_{0}=(x_{0},t_{0})$. Note that $t_{-}<t_{0}<t_{+}$. We now use the trajectories constructed in Section 2 to estimate the right hand side.
Using the chain rule, (1.1), (2.5) with the choice (2.13) we get
(3.2) ${\mathcal{I}}^{\pm}$ =±$=\pm\fint_{Q^{0}}\Big{(}f(z_{\pm})-f(z_{0})\Big{)}\,\varphi(x_{0}){\,\mathrm{d}}z_{0}$
$=\mp\fint_{Q^{0}}\int_{0}^{1}\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}s}f\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$=\mp\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left(\mathcal{T}f\right)\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$\qquad\mp\fint_{Q^{0}}\int_{0}^{1}\left({\frac{\mathrm{d}}{\mathrm{d}s}}X^{(0)}(s)\right)\cdot\left(\nabla_{(0)}f\right)(\Gamma_{\pm}(s))\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
±$\leq\underbrace{\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{\pm}_{1}}$
$\qquad\underbrace{\mp\fint_{Q^{0}}\int_{0}^{1}\sum_{i=0}^{\kappa}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\left(\nabla_{(0)}f\right)(\Gamma_{\pm}(s))\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{\pm}_{2}}.$
Note that the only difference between the two terms ${\mathcal{I}}^{+}$ and ${\mathcal{I}}^{-}$ is the role of $z_{0}$: in the former case it is the past variable, in the latter it is the future variable.
In the following computations to estimate $\mathcal{I}_{1}^{\pm}$ and $\mathcal{I}_{2}^{\pm}$, the local and the non-local case can be treated similarly, upon replacing the integration domain over $Q^{0}$ by ${\mathbb{R}}^{d_{0}}\times Q^{0}_{v}$ in the non-local case, where $Q^{0}_{v}\subset{\mathbb{R}}^{N+1-d_{0}}$ is the domain of the cylinder $Q^{0}$ for the variables $(x^{(1)},\dots,x^{(\kappa)},t)$, that is we slice out the $x^{(0)}$ variable.
We now integrate by parts the terms ${\mathcal{I}}^{\pm}_{1}$ after a change of variables $x_{0}\mapsto y:=\Phi^{s}_{\pm}(x_{0})$ for $s,t_{0}$ fixed ($\Phi^{s}_{\pm}$ is the affine map defined in (2.11), which is invertible for $s\not=0$):
${\mathcal{I}}_{1}^{\pm}$ =±$=\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
=±$=\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(\Phi^{s}_{\pm}(x_{0}),st_{0}+(1-s)t_{\pm}\right)\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\mathcal{A}f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\,\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=\Phi^{s}_{\pm}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then use the first bound in (1.13) on the operator $\mathcal{A}$ in assumption (H):
|$\left|\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}\right|$ $\lesssim_{\varphi}\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{\beta}$
$\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-1-\alpha_{i}}\right)^{\beta}\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-\beta-\beta\alpha_{i}}\right)$
and deduce finally
${\mathcal{I}}_{1}^{\pm}\lesssim|\delta_{\pm}|\fint_{Q^{0}}\int_{0}^{1}\left|\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\right|\left(\sum_{i=0}^{\kappa}s^{-\beta-\beta\alpha_{i}}\right){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
Now let us turn to ${\mathcal{I}}_{2}^{\pm}$ in (3.2). We find, using the same change of variables,
${\mathcal{I}}_{2}^{\pm}$ $=\mp\int_{0}^{1}\fint_{Q^{0}}\sum_{i=0}^{\kappa}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\nabla_{(0)}f(\Gamma_{\pm}(s))\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}$
$=\mp\frac{1}{|Q^{0}|}\sum_{i=0}^{\kappa}\int_{0}^{1}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\nabla_{(0)}f\left(y,st_{0}+(1-s)t_{\pm}\right)$
×$\hskip 227.62204pt\times\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\mp\frac{1}{|Q^{0}|}\sum_{i=0}^{\kappa}\int_{0}^{1}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\left(y,st_{0}+(1-s)t_{\pm}\right)$
×$\hskip 170.71652pt\times(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\mp\frac{1}{|Q^{0}|}\sum_{i=0}^{\kappa}\int_{0}^{1}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)$
×$\hskip 99.58464pt\times\left\{(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then observe that
$\left|(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right|\lesssim\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{1-\beta}\lesssim\left(\sum_{j=0}^{\kappa}s^{-1-\alpha_{j}}\right)^{1-\beta}$
and we deduce
${\mathcal{I}}_{2}^{\pm}\leq\frac{1}{|Q^{0}|}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+\beta(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
It follows from (3.1) and (3.2) that we are left with estimating
${\mathcal{J}}:=\fint_{Q^{\pm}}\big{\{}{\mathcal{I}}^{\pm}_{1}\big{\}}_{+}{\,\mathrm{d}}z_{\pm}+\fint_{Q^{\pm}}\big{\{}{\mathcal{I}}^{\pm}_{2}\big{\}}_{+}{\,\mathrm{d}}z_{\pm}.$
The previous estimates imply
(3.3) ${\mathcal{J}}$ $\lesssim\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i=0}^{\kappa}\frac{1}{s^{(\alpha_{i}+1)\beta}}\right)\left|\left({\mathcal{A}}f\right)(\Gamma_{\pm}(s))\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$+\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+\beta(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}.$
In order to balance optimally the singularities we then choose the $\alpha_{i}$’s all close to $(2+\beta)^{-1}-1\in(-1,0)$ pairwise distinct.
Remark 8.
Note also that, by adopting the mollified formulation (1.19) and by keeping the test functions $\psi_{\pm}$ until the integral ${\mathcal{J}}$ above, one could use the variable $z_{0}$ to integrate the test function in the non-singular region $s\in[s_{0},1]$, keep an $L^{\infty}$ control of the test function in the singular region, and optimise $s_{0}$ in order to deduce a gain of integrability.
Then (3.3) simplifies into, for some $\varepsilon$ as small as wanted,
${\mathcal{J}}$ $\lesssim\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}s^{-\frac{\beta}{2+\beta}-\varepsilon}\left|\left({\mathcal{A}}f\right)(\Gamma_{\pm}(s))\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$+\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}s^{-1+\frac{\beta}{2+\beta}-\varepsilon}\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}.$
We are now in a position to use not only the intermediate variable $z_{0}$, but also the future/past variables $z_{\pm}$ for a change of variables, since we are now integrating in both $Q^{0}$ and $Q_{\pm}$. Note that it was not possible to use the integration in $Q_{\pm}$ before because of the positive value around the $Q^{0}$ integral.
We split ${\mathcal{J}}$ as follows, given $s_{0}\in(0,1)$,
$\int_{Q^{\pm}}\int_{Q^{0}}$ $\int_{0}^{1}\Big{(}\cdots\Big{)}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}{\,\mathrm{d}}z_{\pm}$
$=\underbrace{\int_{Q^{\pm}}\int_{Q^{0}}\int_{0}^{s_{0}}\Big{(}\cdots\Big{)}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}{\,\mathrm{d}}z_{\pm}}_{=:{\mathcal{J}}^{\pm}_{1}}+\underbrace{\int_{Q^{\pm}}\int_{Q^{0}}\int_{s_{0}}^{1}\Big{(}\cdots\Big{)}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}{\,\mathrm{d}}z_{\pm}}_{=:{\mathcal{J}}^{\pm}_{2}}.$
The two changes of variables on each part are represented in Figure 4.
[…]
Figure 4. The change of variables that we use in the proof for some fixed $s\in(0,1)$. For $s\in(0,s_{0})$ we use $\Psi_{\pm}$ that map $z_{\pm}$ onto $\Gamma_{\pm}$, whereas for $s\in(s_{0},1)$ we use $\Phi_{\pm}$ that map $z_{0}$ onto $\Gamma_{\pm}$.
The part ${\mathcal{J}}^{\pm}_{2}$ is controlled by using the same change of variables $\Phi^{s}_{\pm}$ as before:
${\mathcal{J}}^{\pm}_{2}$ $=\int_{Q^{0}}\int_{s_{0}}^{1}s^{-1+\frac{\beta}{2+\beta}-\varepsilon}\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(y,st_{0}+(1-s)t_{\pm}\right)\right|\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$+\int_{Q^{0}}\int_{s_{0}}^{1}s^{-\frac{\beta}{2+\beta}-\varepsilon}\left|{\mathcal{A}}f\left(y,st_{0}+(1-s)t_{\pm}\right)\right|\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$\lesssim_{s_{0}}\int_{\tilde{\Omega}}\left|(-\Delta_{(0)})^{\frac{\beta}{2}}f(z)\right|{\,\mathrm{d}}z+\int_{\tilde{\Omega}}\left|{\mathcal{A}}f(z)\right|{\,\mathrm{d}}z\lesssim\int_{\tilde{\Omega}}\left|{\mathcal{A}}f(z)\right|{\,\mathrm{d}}z,$
where we have used first that the integral in $s$ avoids the singularity at 00, and second assumption (1.13) in (H). To control the part ${\mathcal{J}}^{\pm}_{1}$ we parametrise $\Gamma_{\pm}(s)$ by the $z_{\pm}$ coordinates, for $s\in[0,s_{0}]$ and $t_{0},t_{\pm}$ all fixed:
$(x_{\pm},t_{\pm})\to\Gamma_{\pm}(s)=(\bar{\Gamma}^{s}_{\pm},t_{0}s+(1-s)t_{\pm})=:(\Psi^{s}_{\pm}(x_{\pm}),t_{0}s+(1-s)t_{\pm})$
Since $s$ does not approach $1$, we can prove that this change of variables in not singular. The mappings $\Psi_{\pm}^{s}$ are determined by solving (2.11) for $x_{\pm}$ instead of $x_{0}$. It yields $\Psi_{\pm}^{s}x_{\pm}=\mathfrak{a}_{\pm}^{s}x_{\pm}+\mathfrak{b}_{\pm}^{s}$ with the matrix $\mathfrak{a}_{\pm}^{s}$ and vector $\mathfrak{b}_{\pm}^{s}$ given by
$\mathfrak{a}^{s}_{\pm}:=\left({\mathsf{T}}_{\pm}(s)-{\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}{\mathsf{T}}_{\pm}(1)\right)\quad\mathfrak{b}^{s}_{\pm}:={\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}x_{0}$
which depend only on $s$ and $x_{0}$. Since we have proved in Lemma 7 that ${\mathsf{W}}(s)={\mathcal{O}}\left(s^{d_{0}\frac{(\kappa+1)(\kappa+2)}{2}+d_{0}\sum_{i=0}^{\kappa}\alpha_{i}}\right)$ it follows that ${\mathsf{W}}^{\delta}_{\pm}(s)={\mathcal{O}}\left(s^{d_{0}\frac{(\kappa+1)(\kappa+2)}{2}+d_{0}\sum_{i=0}^{\kappa}\alpha_{i}}\right)$ and thus ${\mathsf{W}}^{\delta}_{\pm}(s)$ goes to zero as $s\to 0$. Since ${\mathsf{T}}^{\delta}_{\pm}$ is a lower triangular matrix with unitary diagonal it is invertible for all $s$, and finally $\mathfrak{a}^{s}_{\pm}$ is invertible with uniformly bounded inverse on $s\in[0,s_{0}]$ for $s_{0}>0$ small enough. We apply this change of variables to estimate ${\mathcal{J}}_{1}^{\pm}$:
${\mathcal{J}}^{\pm}_{1}$ $=\int_{Q^{0}}\int_{0}^{s_{0}}s^{-\frac{\beta}{2+\beta}-\varepsilon}\left|\left({\mathcal{A}}f\right)\left(y,st_{0}+(1-s)t_{\pm}\right)\right|\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{a}^{s}\right|}$
$+\int_{Q^{0}}\int_{0}^{s_{0}}s^{-1+\frac{\beta}{2+\beta}-\varepsilon}\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(y,st_{0}+(1-s)t_{\pm}\right)\right|\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{a}^{s}\right|}$
$\lesssim_{s_{0}}\int_{\tilde{\Omega}}\left|{\mathcal{A}}f(z)\right|{\,\mathrm{d}}z$
which concludes the proof. Note that in the local case $\tilde{\Omega}$ coincides with the solution domain $\Omega$, whereas in the non-local case, the domain for the first $d_{0}$ variables of $\tilde{\Omega}$ coincides with the whole space ${\mathbb{R}}^{d_{0}}$, so that $\tilde{\Omega}={\mathbb{R}}^{d_{0}}\times B_{R}\times\dots\times B_{R}\times(-5,0)$.
4. Proof of the Harnack inequality
In this section we briefly explain how to adapt the approach of [8] for deducing the intermediate value of De Giorgi from the Poincaré inequality we have proved. This approach was developed in the local case with one commutator and with a weaker version of the Poincaré inequality. A non-local extension, with one commutator, was developed in [13]. This approach can be described as a “trajectory viewpoint” on the De Giorgi theory, and it is fully quantitative. The first De Giorgi lemma writes in its simplest form
Lemma 9 (First Lemma of De Giorgi).
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
  • (i)
    ${\mathcal{A}}:=\sqrt{A}\nabla_{(0)}$ with $A\in[\Lambda^{-1},\Lambda]$ measurable symmetric matrix (local case),
  • (ii)
    ${\mathcal{A}}=\sqrt{a}(-\Delta_{v})^{\beta/2}$ with $a\in[\Lambda^{-1},\Lambda]$ scalar measurable (non-local case).
Let $f$ be a sub-solution to (1.1) on $Q_{1}$ with $0\leq f$ in $Q_{1}$ (local case) or on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-1,0]$ with $0\leq f\leq 1$ (non-local case). Then there exist $\varepsilon>0$ and $r\in(0,1)$ so that $\int_{Q_{1}}f^{2}\leq\varepsilon$ implies $f\leq 1/2$ on $Q_{r}$.
Note that in the local case, a stronger statement holds where the $L^{\infty}$ norm of $f$ on $Q_{r}$ is controlled by a constant (depending on ${\mathcal{A}}$, ${\mathcal{B}}$ and $r$) times the $L^{2}$ norm of $f$ on $Q_{1}$. In the non-local case, variants of this stronger statement exist if $f$ is not only a sub-solution but also a super-solution; we refer to [12, 14]. The proof is based on an energy estimate and the gain of integrability of the fundamental solution of the Kolmogorov equation, local or non-local. This is classical and we refer to [8, 13, 2, 19, 10, 1]. The non-local case with more than one commutator is obtained by scaling arguments or by calculating the fundamental solutions in Fourier.
The second De Giorgi Lemma, also known as the intermediate value Theorem, writes
Lemma 10 (Second Lemma of De Giorgi).
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
  • (i)
    ${\mathcal{A}}:=\sqrt{A}\nabla_{(0)}$ with $A\in[\Lambda^{-1},\Lambda]$ measurable symmetric matrix (local case),
  • (ii)
    ${\mathcal{A}}=\sqrt{a}(-\Delta_{v})^{\beta/2}$ with $a\in[\Lambda^{-1},\Lambda]$ scalar measurable (non-local case).
Let $f$ be a sub-solution to (1.1) with $f\leq 1$ on $Q_{1}$ (local case) or $0\leq f\leq 1$ on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-3,0]$ (non-local case). Let $\delta_{1},\delta_{2}\in(0,1)$. Then there are constants $r_{0}>0$ and $\nu\in(0,1)$ and $\theta\in(0,1)$ such that whenever
|$|\{f\leq 0\}\cap Q_{r_{0}}^{-}|\geq\delta_{1}|Q_{r_{0}}^{-}|\qquad\textrm{ and}\qquad|\{f\geq 1-\theta\}\cap Q_{r_{0}}|\geq\delta_{2}|Q_{r_{0}}|,$
there holds
|$\big{|}\{0<f<1-\theta\}\cap Q_{1/2}\big{|}\geq\nu|Q_{1/2}|$
in the local case, or
|$\big{|}\{0<f<1-\theta\}\cap B_{(\frac{1}{2})^{1+\kappa\beta}}\times\cdots\times B_{(\frac{1}{2})^{1+\beta}}\times[-3,0]\big{|}\geq\nu|Q_{1/2}|$
in the non-local case.
The proof that this lemma follows from our Poincaré inequality and the first lemma of De Giorgi by using the argument of [8, Subsection 3.2] in the local case (in fact the proof is simpler since we do not have the error terms in the Poincaré inequality that were present in [8]), and it follows from [13, Subsection 5.2] in the non-local case (note that we have removed the mollifications $F_{i}$ of the cutoff functions in the statement of [13] since they can be used in the proof but removed from the statement straightforwardly by making the border region small enough). These proofs are done in the case of one commutator but they extend straightforwardly to the case of several commutators because they do not depend on the transport part ${\mathcal{T}}=\partial_{t}+{\mathcal{B}}$ of the equation, but only on ${\mathcal{A}}$ and the Poincaré inequality and first lemma of De Giorgi.
With the first and second De Giorgi lemma at hand, one can deduce the “measure-to-pointwise estimate” and the weak Harnack inequality by following [8] in the local case, and [13] in the non-local case. These proofs are independent of the equation and only depend on the previously established functional inequalities. The only minor change is that the kinetic cylinders have a slightly more complicated scaling on the variables $(x^{(\kappa)},\dots,x^{(0)},t)$ but the technical changes needed are, although slightly tedious, clear.